rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#10626 — FS#14493 — bhs2-17b-n56

Attached to Project— Network
Maintenance
Beauharnois → BHS-2
In progress
0%
In order to correct a dysfunction associated with port-security, we must restart the switch.
During the operation, the traffic will be supported by the other switch of the pair (bhs2-17a-n56).

This reboto is expecyed on Thursday, August 27th from 10am CEST (4am EST).
Comment by OVH - Thursday, 27 August 2015, 12:54PM

We've isolated the switch. We will start the reboot.